OpenStudy (anonymous):

Factor completely 64x2 – 1

OpenStudy (anonymous):

(8x-1) (8x +1)

OpenStudy (anonymous):

64x^2 = 1 x^2 = 1/64 x = +/- sqrt(64) x = 1/8 and -1/8 (x+1/8)(x-1/8) ;)